OpenStudy (anonymous):

What is the y-intercept of the line given by the equation below? Enter your answer as an ordered pair and enter any fractions using the slash mark ( / ) for the fraction bar. y = x +

OpenStudy (texaschic101):

I do not understand your question. What is the entire equation of the line given ?

OpenStudy (anonymous):

y = x +

OpenStudy (anonymous):

y=7/4+1/3

OpenStudy (texaschic101):

ok...in y = mx + b form, the y intercept is in the b position. So in y = 7/4 + 1/3....the y intercept is 1/3

OpenStudy (texaschic101):

wait..is that supposed to be 7/4x ?

OpenStudy (anonymous):

no you were right the first time

OpenStudy (texaschic101):

another way to find the y intercept is to sub in 0 for x and solve for y